# Gravitational Wave Signatures of Primordial Black Hole Reheating in Upcoming Interferometry Missions

### v1

@@ 4.2.1 Statistical analysis in the light of the detectors
On the other hand, for LISA (see, for instance, Fig. 12(b) ), the parameter span starts from $M_{\rm in}\sim 2\times 10^{4}$ g. For PBH masses below this threshold, the SNR remains $<1$ for all possible ranges of $w_{\phi}$ and $\beta$. Furthermore, for LISA, the formation mass consistent with $\rm SNR>1$ lies in the range $M_{\rm in}\sim(2\times 10^{4}-5\times 10^{8})$ g, where the upper limit corresponds to the maximum mass value that evaporated just before BBN.
@@
Notably, for both ET and LISA, in the high SNR region (i.e. SNR $>1$), we find consistently low uncertainties (i.e. $\Delta w_{\phi}/w_{\phi},\Delta\beta/\beta<1$), reinforcing the idea that induced GWs from the adiabatic source can be measured with high precision. Thus, in agreement with the SNR analysis, the Fisher analysis further confirms that both LISA and ET are well-suited for detecting induced GWs from the adiabatic source with minimal uncertainties.
@@ 5 Combined analysis with primordial source and PBH domination
The two sources of induced GWs are independent and contribute to the spectrum of the induced GWs simultaneously in the PBH reheating scenario. Consequently, the most realistic approach is to assess the detectability of the combined spectrum at the detectors. The resulting present-day energy density spectrum of the induced GWs can be expressed as
@@ 5.1 Statistical analysis in the light of the detectors
Similarly to the earlier study, we have estimated the SNR for ET and LISA in Figs. 18 and 19, respectively, for two representative values of $M_{\rm in}$ ($10^{5}$ g and $10^{7}$ g). The other annotations and shaded regions have the same physical interpretation as before. Fig. 17 suggests that the peaks in the spectrum, originating from two different sources, dominate each other for different PBH parameters. Consequently, some of the parameter space may be reduced due to the overproduction of GWs compared to individual sources. This effect is evident by comparing the adiabatic scenario (Fig. 11(a)) with the combined scenario (Fig. 19(a)), where the SNR $>1$ region in the $w_{\phi}-\beta$ plane shrinks in the combined case. To illustrate this effect, we have explicitly compared the GW overproduction limit of the combined spectrum with those from individual sources for different masses in Fig. 20, showing that the combined scenario follows a tighter limit than the individual sources.
@@
Based on the basic setup described in Sec. 3.2, we have conducted the MCMC analysis for ET and LISA, considering the combined scenario, resulting in an estimation of the parameters under consideration using mock data generated from the instrumental specifications and possible sources of noise for both the missions. The priors adopted for the parameters are listed in Table-2. To generate the mock data, we have considered the fiducial parameter set by ($w_{\phi},{\rm log}_{10}(\beta),{\rm log}_{10}(\frac{M_{\rm in}}{1{\rm g}})$) $\equiv$ ($0.5,-9,5$) for both detectors. This choice ensures that resulting GW spectrum satisfies SNR $>1$ and relative uncertainties $<1$, for both detectors. Finally, the mock catalogue itself is generated following methodology outlined in Sec. 3.2. Using on the mock catalogues, we have presented 1- and 2-dimensional posterior distributions for ET and LISA in Fig. 27, with the corresponding statistical results summarized in Table-4.
@@ 6 Summary and outlook
[…]
@@
The analysis presented in the article can be extended further in a number of directions. This present analysis focuses only on Schwarzschild PBHs with a monochromatic mass distribution. A key extension would be to explore spinning PBHs and an extended mass function, as both could significantly impact the induced GW spectrum. Understanding how an extended mass distribution influences the amplitude, peak frequency, and shape of the GW signal will be crucial for refining detection prospects. Additionally, our model-independent approach to PBH parameters can be further developed by linking constraints to specific formation mechanisms. Connecting induced GWs to scenarios like inflationary dynamics, cosmic strings, or phase transitions could provide deeper insights into early Universe physics.
@@ Appendix B Calculation of cutoff function
[…]
@@
[…]
@@
Acknowledgements
Authors sincerely thank Jan Tränkle and Guillem Doménech for their fruitful discussion. Authors also acknowledge the use of publicly available codes emcee [108] and GetDist [109], and thank the computational facilities of Technology Innovation Hub, ISI Kolkata and of the Pegasus cluster of the high performance computing (HPC) facility at IUCAA, Pune. DP thanks ISI Kolkata for financial support through Senior Research Fellowship. MRH acknowledges ISI Kolkata for providing financial support through Research Associateship. SP thanks the ANRF, Govt. of India for partial support through Project No. CRG/2023/003984.

### v2

@@ 4.2.1 Statistical analysis in the light of the detectors
On the other hand, for LISA (see, for instance, Fig. 12(b) ), the parameter span starts from $M_{\rm in}\sim 2\times 10^{4}$ g. For PBH masses below this threshold, the SNR remains <1<1 for all possible ranges of $w_{\phi}$ and $\beta$. Furthermore, for LISA, the formation mass consistent with $\rm SNR>1$ lies in the range $M_{\rm in}\sim(2\times 10^{4}-5\times 10^{8})$ g, where the upper limit corresponds to the maximum mass value that evaporated just before BBN.
@@
Notably, for both ET and LISA, in the high SNR region (i.e. SNR >1>1), we find consistently low uncertainties (i.e. $\Delta w_{\phi}/w_{\phi},\Delta\beta/\beta<1$), reinforcing the idea that induced GWs from the adiabatic source can be measured with high precision. Thus, in agreement with the SNR analysis, the Fisher analysis further confirms that both LISA and ET are well-suited for detecting induced GWs from the adiabatic source with minimal uncertainties.
@@
Figure 16: Figure is same as Fig. 13, but for relative uncertainties on $\beta$, considering LISA.
@@ 5 Combined analysis with primordial source and PBH domination
The two sources of induced GWs are independent and contribute to the spectrum of the induced GWs simultaneously in the PBH reheating scenario. Consequently, the most realistic approach is to assess the detectability of the combined spectrum at the detectors. The resulting present-day energy density spectrum of the induced GWs  [^7] Note that our analysis is carried out within the framework of linear cosmological perturbation theory. can be expressed as
@@ 5.1 Statistical analysis in the light of the detectors
Similarly to the earlier study, we have estimated the SNR for ET and LISA in Figs. 18 and 19, respectively, for two representative values of $M_{\rm in}$ ($10^{5}$ g and $10^{7}$ g). The other annotations and shaded regions have the same physical interpretation as before. Fig. 17 suggests that the peaks in the spectrum, originating from two different sources, dominate each other for different PBH parameters. Consequently, some of the parameter space may be reduced due to the overproduction of GWs compared to individual sources. This effect is evident by comparing the adiabatic scenario (Fig. 11(a)) with the combined scenario (Fig. 19(a)), where the SNR >1>1 region in the $w_{\phi}-\beta$ plane shrinks in the combined case. To illustrate this effect, we have explicitly compared the GW overproduction limit of the combined spectrum with those from individual sources for different masses in Fig. 20, showing that the combined scenario follows a tighter limit than the individual sources.
@@
Based on the basic setup described in Sec. 3.2, we have conducted the MCMC analysis for ET and LISA, considering the combined scenario, resulting in an estimation of the parameters under consideration using mock data generated from the instrumental specifications and possible sources of noise for both the missions. The priors adopted for the parameters are listed in Table-2. To generate the mock data, we have considered the fiducial parameter set by ($w_{\phi},{\rm log}_{10}(\beta),{\rm log}_{10}(\frac{M_{\rm in}}{1{\rm g}})$) $\equiv$ (0.5,9,50.5,-9,5) for both detectors. This choice ensures that resulting GW spectrum satisfies SNR >1>1 and relative uncertainties <1<1, for both detectors. Finally, the mock catalogue itself is generated following methodology outlined in Sec. 3.2. Using on the mock catalogues, we have presented 1- and 2-dimensional posterior distributions for ET and LISA in Fig. 27, with the corresponding statistical results summarized in Table-4.
@@ 6 Summary and outlook
  • Isocurvature fluctuations: Our analysis reveals distinct detection capabilities for two different interferometry missions, ET and LISA. While ET demonstrates strong sensitivity across a broad PBH mass range, $M_{\rm in}\in(0.5-3.34\times 10^{7})$ g, LISA is only sensitive to PBH masses above $10^{8}$ g. However, ET is not expected to detect the spectrum for PBHs with a formation mass exceeding $4\times 10^{7}$ g, as the formation peak lies beyond its sensitivity range, highlighting a complementary detection window between LISA and ET. The parameter space in the $\beta-w_{\phi}$ plane is explored, constrained by the requirement $\text{SNR}>1$ and limits from GW overproduction (i.e., the $\Delta N_{\rm eff}$ bound at BBN). One notable outcome is that the detectable region $(\text{SNR}>1$) in the $\beta-w_{\phi}$ plane for ET is significantly large for a PBH formation mass around $M_{\rm in}\sim 10^{6}$ g. This is expected, as the peak frequency, associated with the UV cutoff scale $k_{\rm UV}$ for the masses around $10^{6}$ g, falls within the range where the noise spectrum $\Omega_{\rm GW}^{\rm noise}$ for ET attains its minimum. In contrast, for LISA, achieving $\text{SNR}>1$ proves highly challenging, making it less favorable for detecting the isocurvature source of induced GWs.
@@
The analysis presented in the article can be extended further in a number of directions. This study investigates Schwarzschild primordial black holes with a monochromatic mass distribution. A significant extension would involve analyzing spinning PBHs and adopting an extended mass function, as these could profoundly influence the induced gravitational wave spectrum. An extended mass distribution typically prolongs the transition from matter to radiation domination, leading to a pronounced suppression of the GW signal. This effect arises from a weakened resonance mechanism, and subtle cancellation further suppresses the signal. Prior studies, such as those in references [118, 119, 120, 67, 63], demonstrate that extended mass distributions broaden evaporation timescales, resulting in a smoother transition and a suppression of the induced gravitational wave signal by an order of magnitude. Additionally, our model-independent approach to PBH parameters can be further developed by linking constraints to specific formation mechanisms. Connecting induced GWs to scenarios like inflationary dynamics, cosmic strings, or phase transitions could provide deeper insights into early Universe physics.
At this point, we emphasize that our analysis is based on linear cosmological perturbation theory, which offers a clear and tractable framework for predicting the gravitational wave spectrum, including the characteristic double-peak feature. However, if the early matter-dominated phase is sufficiently long, small-scale density perturbations can enter the non-linear regime, as their amplitude grows with the scale factor. As a result, adopting a fixed cutoff at $k_{\rm UV}$ may, in some cases, lead to an overestimation. To accurately capture the power spectrum in the non-linear regime, one must resort to detailed numerical simulations [121] or analytical techniques such as kinetic field theory [122, 123].
@@
Acknowledgements
Authors sincerely thank Jan Tränkle and Guillem Doménech for their fruitful discussion, and the anonymous referee for valuable insights and suggestions in the manuscript. Authors also acknowledge the use of publicly available codes emcee [111] and GetDist [112], and thank the computational facilities of Technology Innovation Hub, ISI Kolkata and of the Pegasus cluster of the high performance computing (HPC) facility at IUCAA, Pune. DP thanks ISI Kolkata for financial support through Senior Research Fellowship. MRH acknowledges ISI Kolkata for providing financial support through Research Associateship. SP thanks the ANRF, Govt. of India for partial support through Project No. CRG/2023/003984.
@@ Appendix B Calculation of cutoff function
$\Theta^{\rm(iso)}_{\rm UV}(k)$ $\equiv$ $\int_{-s_{0}(k)}^{s_{0}(k)}ds\frac{\left(s^{2}-1\right)^{2}}{\left(1-c_{s}^{2}s^{2}\right)^{5/3}}$
@@
$\Theta^{\rm(ad)}_{\rm UV}(k)\equiv$ $\int_{-s_{0}(k)}^{s_{0}(k)}ds\left(s^{2}-1\right)^{2}\left(1-c_{s}^{2}s^{2}\right)^{n_{\rm eff}(w_{\phi})}$ (B.2)